15x - 9y = 0
2x + 3y = 21

Answer by tommyt3rd(5050)

from 1)
y=5/3 x

from 2)
2x+3(5/3 x) = 21
2x+5x=21
7x=21
x=3

then
y=5/3 (3)
y=5
solution:
(3,5)